OpenStudy (anonymous):

Anyone know what -15-20 is? I just need to double check...no calculator. :0(

OpenStudy (saifoo.khan):

-35

OpenStudy (anonymous):

-35

OpenStudy (anonymous):

Thank you!

OpenStudy (shadowfiend):

The easiest way to do this is to pull the negative out. -15 - 20 = -(15 + 20) So, you can add 15 and 20, get 35, and then you still have the negative sign outside.

OpenStudy (anonymous):

And 12-(-4) = 12 + 4 = 16, right? or wrong?